OpenStudy (anonymous):

Solve by substitution: y = 2x - 3 y = -x + 1 Please help, I really don't get how to do this

OpenStudy (anonymous):

@tHe_FiZiCx99

OpenStudy (the_fizicx99):

K, to substitute plug in y into the other equation, y = 2x - 3 y = -x + 1 -x + 1 = 2x - 3 Now can you solve that?

OpenStudy (anonymous):

I'm still not sure, i know how to substitute but this question has just been confusing me :/ @tHe_FiZiCx99

OpenStudy (the_fizicx99):

The term itself says it :P Substitute in the " y = " equation into the other, I subbed the second equation into the first,

OpenStudy (anonymous):

solve for x after you substitute

OpenStudy (the_fizicx99):

I got dis >.<

OpenStudy (anonymous):

how can i solve for x though?

OpenStudy (the_fizicx99):

-x + 1 = 2x - 3 Add both sides x -x + 1 = 2x - 3 +x +x 1 = 3x - 3 Add both sides 3 to leave x alone, 1 = 3x -3 +3 +3 4 = 3x Divide to isolate x x = 4/3 NOW, to find Y plug in x into either equation and solve.

OpenStudy (anonymous):

ok, whenever people say "solve for x" you just move the number around when problems are like this: -x + 1 = 2x - 3 - 1= - 1 -------------- -x = 2x -4 -2x=-2x ------------ -3x= -4 and so on, get it?

OpenStudy (anonymous):

it does not matter which number you start first from either side. get me?

OpenStudy (anonymous):

I'm still a little confused but I'm catching on

OpenStudy (the_fizicx99):

Ok, so now plug in 4/3 to find y, y = 2(4/3) - 3 First multiply 4/3 * 2

OpenStudy (anonymous):

so y = -0.34

OpenStudy (the_fizicx99):

y = -1/3 :p -1/3 = -0.33333333333333333333333333333333

OpenStudy (the_fizicx99):

4/3 * 2 = 2.6666666666666666666666666666667 2.6666666666666666666666666666667 - 3 = -0.33333333333333333333333333333333

OpenStudy (the_fizicx99):

You plug both y and x to see if its a true statement, -1/3 = 2(4/3) - 3 -1/3 = -1/3 :P

OpenStudy (anonymous):

-1.3 is y

OpenStudy (anonymous):

and 2.6 is x

OpenStudy (anonymous):

@tHe_FiZiCx99

OpenStudy (anonymous):

never mind, -0.3 is y and -1.3 is x right?

OpenStudy (the_fizicx99):

No, you're not solving for them, your checking if they give you a true statement. x = 4/3 y = -1/3

OpenStudy (anonymous):

oh, ok thank you soooo much!!

OpenStudy (the_fizicx99):

yw